OpenStudy (kidthatbro8):

MEDAL & FAN Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (chosenmatt):

so just do what it says

OpenStudy (anonymous):

First find the slope. We have two points: (3, 285) and (6, 510) Plug them into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{510-285}{6-3}\)

OpenStudy (anonymous):

Can you simplify that?

OpenStudy (chosenmatt):

yeah

OpenStudy (anonymous):

@Kidthatbro8

OpenStudy (kidthatbro8):

225 / 3 ?

OpenStudy (anonymous):

Yep, now divide that.

OpenStudy (kidthatbro8):

75 ?

OpenStudy (anonymous):

Yes, so our slope is 75. Now we can plug that in along with one of our given points into point-slope form.

OpenStudy (anonymous):

\(y - y_1 = m(x - x_1)\) Where \(x_1\) is our x-value in our point, \(y_1\) is the y-value in our point, and \(m\) is our slope. Let's use the point \((3, 285)\). So in this case: \(x_1 = 3\) \(y_1 = 285\) \(m = 75\) Plug them in: \(y - 285 = 75(x - 3)\) Now simplify it, distribute 75 into the parenthesis, what do you get?

OpenStudy (kidthatbro8):

i got a very large number. 5400 ?

OpenStudy (anonymous):

How did you get that? Distributing 75 will just multiply it to 'x' and '-3'. \(75 \times x =~?\) \(75 \times -3 =~?\)

OpenStudy (kidthatbro8):

the first equation is a bit confusing, but the second one is -225, right ?

OpenStudy (anonymous):

Yes.. \(75 \times x\) is the same thing as \(75x\).

OpenStudy (anonymous):

So we get: \(y - 285 = 75x - 225\) Now we add 285 to both sides, what's -225 + 285?

OpenStudy (kidthatbro8):

60 ?

OpenStudy (anonymous):

Yep, that gives us: \(y = 75x + 60\) So this is our equation..that takes care of Part A.

OpenStudy (anonymous):

Now for Part B, we change it to function notation, do you know how to do that?

OpenStudy (kidthatbro8):

is that the f(x) thing ?

OpenStudy (anonymous):

Yep, we just replace y with f(x). So that gives us: \(f(x) = 75x + 60\) in function notation..

OpenStudy (anonymous):

That takes care of Part B.

OpenStudy (anonymous):

Now for Part C, you will label the x-axis days, and the y-axis rent. We can just plot the points (3, 285) and (6, 510) then draw a line through it.

OpenStudy (anonymous):

That's all of your answers! :D

OpenStudy (kidthatbro8):

thank you so much ! ;v;

OpenStudy (anonymous):

No problem.